OpenStudy (anonymous):

Which is a solution to the equation 6x + y = –4?

OpenStudy (lgbasallote):

which? where's the list?

OpenStudy (anonymous):

A. (1, 2) B. (1, –10) C. (2, 8) D. (2, –8)

OpenStudy (lgbasallote):

substitute each one to see

OpenStudy (lgbasallote):

is 6(1) + 2 = -4?

OpenStudy (anonymous):

no

OpenStudy (lgbasallote):

therefore cross out A

OpenStudy (anonymous):

ok i see what your saying

OpenStudy (lgbasallote):

6(1) + (-10) = 4?

OpenStudy (lgbasallote):

i mean 6(1) + (-10) = -4?

OpenStudy (anonymous):

no

OpenStudy (lgbasallote):

actually 6 - 10 = -4

OpenStudy (anonymous):

ahhhh

OpenStudy (lgbasallote):

you see now?

OpenStudy (anonymous):

yesss thank you